a chord is 7 cent. from center of circle whose radius is 9 cent in lenght. find the length of chord
Answer by nerdybill(5404)

a chord is 7 cent. from center of circle whose radius is 9 cent in lenght. find the length of chord
.
Draw a diagram of the problem. Label all the information provided. It'll make it easier for you to see the problem.
.
Let x = 1/2 the length of the chord
then applying Pythagorean theorem:
x^2 + 7^2 = 9^2
x^2 + 49 = 81
x^2 = 32
x = sqrt(32)
x = 4sqrt(2)
.
Length of chord:
2x = 2(4sqrt(2)) = 8sqrt(2)